how many grams would be contained in 3.49 x 10^80 molecules of H2SO4?

1
Expert's answer
2022-05-13T05:27:04-0400

Solution:

One mole of any substance contains 6.02×1023 atoms/molecules.

Therefore,

Moles of H2SO4 = (3.49×1080 molecules) × (1 mol / 6.02×1023 molecules) = 5.797×1056 mol H2SO4


The molar mass of H2SO4 is 98.079 g/mol

Therefore,

Mass of H2SO4 = (5.797×1056 mol H2SO4) × (98.079 g H2SO4 / 1 mol H2SO4) = 5.686×1058 g H2SO4

Mass of H2SO4 = 5.686×1058 g


Answer: 5.686×1058 grams of H2SO4